# Further author information: (Send correspondence to Songjiang-Lai.)

### v2

@@ 1 INTRODUCTION
Single Image Super-Resolution (SISR) aims to reconstruct a high-quality image from a low-resolution one. The development of effective super-resolution algorithms has emerged as a pivotal research domain in computer vision owing to its extensive applications. Recent research have integrated the self-attention mechanism into computer vision challenges[1, 2].
@@
CNN-based techniques for Single image Super-Resolution (SISR) have markedly improved the restoration of image texture features. SRCNN[3] was the inaugural model to tackle super-resolution with convolutional neural networks. VDSR[4] implemented residual learning to enhance learning and successfully address the gradient vanishing issue in deep networks. In SRGAN[5], Christian Ledig et al. employed generative adversarial networks to refine super-resolution image production, with the generator converting low-resolution images to high-resolution ones and improving quality via adversarial training. ESRGAN[6] included the Residual Dense Block (RRDB) as a fundamental network component, diminishing perceptual loss by utilizing characteristics prior to activation, hence producing images with more authentic textures. Moreover, researchers persist in suggesting novel structures to retrieve progressively realistic information in super-resolution images. CNN-based networks have demonstrated considerable performance efficacy. Nonetheless, the inductive bias inherent in CNNs constrains the ability of SISR models to capture long-range relationships. The constraints arise from the parameter-dependent scaling of the receptive field and the convolution operator’s kernel size across many layers, potentially neglecting non-local spatial information in pictures.
@@
To enhance the joint modeling of different hierarchical structures in pictures, researchers have taken use of the self-attention mechanism’s benefits in multi-scale processing and long-range dependency modeling. Transformer-based SISR models have been developed to overcome the shortcomings of CNN-based networks by utilizing their capacity to simulate long-range dependencies and improving SISR performance. For instance, super-resolution results have been markedly enhanced by SwinIR[1], which makes use of the Swin Transformer[2]. Furthermore, state-of-the-art results have been produced with a hybrid attention transformer (HAT)[7] that combines an overlapping cross-attention module with window-based self-attention and channel attention.
@@
Despite the successful use of Transformer-based approaches to image restoration problems, there are areas for improvement. Current window-based Transformer networks confine self-attention calculations to a concentrated region. This method evidently results in a constrained receptive field and fails to properly use the feature information from the original picture. This research proposes a hybrid multiaxial aggregation network, termed HAAT, to address the aforementioned issues. HAAT is constructed by integrating Swin-Dense-Residual-Connected Blocks (SDRCB)[8] with Hybrid Grid Attention Blocks (HGAB). HGAB, inspired by GAB[9], integrates channel attention, sparse attention, and window attention, leveraging the global information perception capabilities of channel attention to address the deficiencies of self-attention. Sparse self-attention is used to enhance global feature interactions while maintaining computational efficiency. Meanwhile, to further thrill the potential performance of the model.
@@ 2.1 Swin-Dense-Residual-Connected Block
where []delimited-[][$\cdot$][ ⋅ ] denotes the concatenation of multi-level feature maps produced by the previous layers. $H_{trans}(\cdot)$ refers to the convolution layer with a LeakyReLU activate function for feature transition. The negative slope of LeakyReLU is set to 0.2. Conv1 is the $1\times 1$ convolution layer, which is used to adaptively fuse a range of features with different levels[10]. represents residual scaling factor, which is set to 0.2 for stabilizing the training process[6].
@@ 2.2 Hybrid Grid Attention Block(HGAB)
Figure 2: HGAB Structure.
The structure of Hybrid Grid Attention Block (HGAB) is illustrated in Figure 2. The HGAB consists of a Mix Attention Layer (MAL) and an MLP layer. Regarding the MAL, we first split the input feature $F_{in}$ into two parts by channel: $F_{G}\in R^{(H\times W\times{C/2})}$ and $F_{W}\in R^{(H\times W\times{C/2})}$. Besides, feed the input to another branch to perform channel attention operation. Subsequently, we split $F_{W}$ into two parts by channel again and input them into W-MSA and SW-MSA, respectively. Meanwhile, $F_{G}$ is input into Grid-MSA[9].The computation process of MAL is as follows:
@@
where $X_{W_{1}}$, $X_{W_{2}}$, $X_{G}$ and $X_{C}$ are the output features of W-MSA, SW-MSA, Grid-MSA, and CA, respectively. Furtheremore, it should be noted that we adopt the post-norm method in HGAB to enhance the network training stability. For a given input feature $F_{in}$, the computation process of HGAB is:
@@ 3 EXPERIMENTAL RESULTS
Our HAAT model is trained on DF2K, a substantial aggregated dataset that includes DIV2K[11] and Flickr2K[12]. DIV2K provides 800 images for training, while Flickr2K contributes 2650 images. For the training input, we generate LR versions of these images by applying a bicubic down sampling method with scaling factors of 2, 3, and 4, respectively. To assess the effectiveness of our model, we conduct performance evaluations using well-known SISR benchmark datasets such as Set5[13], Set14[14].
@@
For evaluation, we use all RGB channels and exclude the outermost (2 × scale) border pixels. PSNR and SSIM metrics are employed for assessment. Table 1 shows a quantitative comparison of our method with state-of-the-art approaches such as EDSR[15], RCAN[16], SAN[17], IGN[18], HAN[19], NLSN[20], SwinIR [1], CATA[21], DAT[22] and CDRT[8]. Our method consistently outperforms these methods across all benchmark datasets. Despite having fewer parameters and lower computational requirements, HAAT achieves significantly better results than the other state-of-the-art models.

### v3

@@ 1 INTRODUCTION
Single Image Super-Resolution (SISR) aims to reconstruct a high-quality image from a low-resolution one. The development of effective super-resolution algorithms has emerged as a pivotal research domain in computer vision owing to its extensive applications. Recent research has integrated the self-attention mechanism into computer vision challenges[1, 2].
@@
CNN-based techniques for Single Image Super-Resolution (SISR) have markedly improved the restoration of image texture features. SRCNN[3] was the inaugural model to tackle super-resolution with convolutional neural networks. VDSR[4] implemented residual learning to enhance learning and successfully address the gradient vanishing issue in deep networks. In SRGAN[5], Ledig et al. employed generative adversarial networks to refine super-resolution image generation, with the generator converting low-resolution images to high-resolution ones and improving quality via adversarial training. ESRGAN[6] included the Residual Dense Block (RRDB) as a fundamental network component, diminishing perceptual loss by utilizing characteristics prior to activation, hence producing images with more authentic textures. Moreover, researchers persist in suggesting novel structures to retrieve progressively realistic information in super-resolution images. CNN-based networks have demonstrated considerable performance efficacy. Nonetheless, the inductive bias inherent in CNNs constrains the ability of SISR models to capture long-range relationships. The constraints arise from the parameter-dependent scaling of the receptive field and the convolution operator’s kernel size across many layers, potentially neglecting non-local spatial information in pictures.
@@
To enhance the joint modeling of different hierarchical structures in pictures, researchers have taken advantage of the self-attention mechanism’s benefits in multi-scale processing and long-range dependency modeling. Transformer-based SISR models have been developed to overcome the shortcomings of CNN-based networks by utilizing their capacity to simulate long-range dependencies and improving SISR performance. For instance, super-resolution results have been markedly enhanced by SwinIR[1], which makes use of the Swin Transformer[2]. Furthermore, state-of-the-art results have been produced with a hybrid attention transformer (HAT)[7] that combines an overlapping cross-attention module with window-based self-attention and channel attention.
@@
Despite the successful use of Transformer-based approaches to image restoration problems, there are areas for improvement. Current window-based Transformer networks confine self-attention calculations to a concentrated region. This method results in a constrained receptive field and fails to properly use the feature information from the original picture. This research proposes a hybrid multiaxial aggregation network, termed HAAT, to address these issues. HAAT is constructed by integrating Swin-Dense-Residual-Connected Blocks (SDRCB)[8] with Hybrid Grid Attention Blocks (HGAB). HGAB, inspired by GAB[9], integrates channel attention, sparse attention, and window attention, leveraging the global information perception capabilities of channel attention to address the deficiencies of self-attention. Sparse self-attention is used to enhance global feature interactions while maintaining computational efficiency further enhancing the potential performance of the model.
@@ 2.1 Swin-Dense-Residual-Connected Block
where []delimited-[][$\cdot$][ ⋅ ] denotes the concatenation of multi-level feature maps produced by the previous layers. $H_{trans}(\cdot)$ refers to the convolution layer with a LeakyReLU activation function for feature transition. The negative slope of LeakyReLU is set to 0.2. Conv1 is the $1\times 1$ convolution layer, which is used to adaptively fuse a range of features with different levels[10]. $\alpha$ represents residual scaling factor, which is set to 0.2 for stabilizing the training process[6].
@@ 2.2 Hybrid Grid Attention Block(HGAB)
Figure 2: The structure of HGAB.
The structure of Hybrid Grid Attention Block (HGAB) is illustrated in Figure 2. This advanced hybrid attention block integrates channel, sparse, and self-attention to enhance feature modeling and representation learning beyond traditional hybrid attention approaches. Within the Hybrid Global Attention Block (HGAB), the Multi-Attention Layer (MAL) enables parallel processing of divided input features, boosting efficiency. This design improves global feature interaction and utilizes non-local spatial information, overcoming self-attention limitations while controlling computational complexity. Consequently, it significantly enhances performance in image super-resolution tasks through more sophisticated feature extraction and learning. The HGAB consists of a Mix Attention Layer (MAL) and an MLP layer. Regarding the MAL, we first split the input feature $F_{in}$ into two parts by channel: $F_{G}\in R^{(H\times W\times{C/2})}$ and $F_{W}\in R^{(H\times W\times{C/2})}$. Besides, the input is fed to another branch, where channel attention s applied. Subsequently, we split $F_{W}$ into two parts by channel again and input them into W-MSA and SW-MSA, respectively. Meanwhile, $F_{G}$ is input into Grid-MSA[9].The computation process of MAL is as follows:
@@
where $X_{W_{1}}$, $X_{W_{2}}$, $X_{G}$ and $X_{C}$ are the output features of W-MSA, SW-MSA, Grid-MSA, and CA, respectively. Furthermore, it is worth noting that we adopt the post-norm method in HGAB to enhance the network training stability. For a given input feature $F_{in}$, the computation process of HGAB is as follows:
@@ 3 EXPERIMENTAL RESULTS
Our HAAT model is trained on DF2K, a substantial aggregated dataset that includes DIV2K[11] and Flickr2K[12]. DIV2K provides 800 images for training, while Flickr2K contributes 2650 images. For the training input, we generate LR versions of these images by applying the bicubic down sampling method with scaling factors of 2, 3, and 4. To assess the effectiveness of our model, we conduct performance evaluations using well-known SISR benchmark datasets such as Set5[13] and Set14[14].
@@
For evaluation, we use all RGB channels and exclude the outermost (2 × scale) border pixels. PSNR and SSIM metrics are employed for assessment. Table 1 shows a quantitative comparison of our method with state-of-the-art approaches such as EDSR[15], RCAN[16], SAN[17], IGN[18], HAN[19], NLSN[20], SwinIR [1], CATA[21], DAT[22] and CDRT[8]. Our method consistently outperforms these methods across all benchmark datasets. Therefore, it can be clearly summarized HAAT achieves significantly better results than the other state-of-the-art models.